I have these questions below that I feel would have the same type of set up, however, I feel like I do not know where to start. I was trying to help someone and I did not know where to begin. Please help!!! And thank you in advance!
If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working togehter to mix the 20 drinks?
Jim can fill apool carrying buckets of water in 30 minutes. Sue can do the same job in 45 minutes. Tony can do the same job in 1 1/2 hours. How quickly can all three fill the pool together?
If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa and Tom work together to complete it?
Two cyclists start biking from a trail's start 3 hours apart. The second cyclist travels at 10 miles per hour and starts 3 hours after teh first cyclist who is traveling at 6 miles per hour. How much time will pass before the second cyclist catches up with the first from the time the second cyclist started biking?

Answer by oscargut(891)

If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working togehter to mix the 20 drinks
Steven can mix 20 drinks in 5 minutes
then Steven can mix 20/5 = 4 drinks per minute
Sue can mix 20 drinks in 10 minutes
then Sue can mix 20/10 = 2 drinks per minute
Jack can mix 20 drinks in 15 minutes
then Jack can mix 20/15 = 4/3 drinks per minute
So working together they can mix 4+2+4/3 drinks per minute
They do 22/3 drinks per minute
They do 20 drinks in 20/(22/3)=60/22= 30/11 minutes
Answer: they can mix 20, drinks in 30/11 minutes (2.72 minutes aprox)